### v1

@@ 4.1 Large Archive Preserves Non-dominated Solutions
Throughout the process of SPEA2, the next population $P$ consists of the mutated individuals from the archive $A$. That is, the archive $A$ in SPEA2 actually plays the role of population, while the population $P$ acts as offspring. We now show that when the archive size $\bar{\mu}$ is always no less than the maximum cardinality of the set of non-dominated solutions having different objective vectors for any $m$-objective function $\bm{f}$, SPEA2 will preserve the non-dominated set on $\bm{f}$.
@@
For SPEA2 solving $\bm{f}$, let $S$ denote the non-dominated set of solutions in the union of the population $P$ and archive $A$. If the size $\bar{\mu}$ of archive is always no less than $|\bm{f}(S)|$, then SPEA2 preserves the non-dominated set.
@@
Consider that all the non-dominated solutions in $P\cup A$ have been added to $A^{\prime}$ (line 4 of Algorithm 1). If the size of $A^{\prime}$ is at most $\bar{\mu}$, then all the non-dominated solutions will survive to the next generation. Thus, the lemma holds. If the size of $A^{\prime}$ is larger than $\bar{\mu}$ (i.e., $|S|>\bar{\mu}$), SPEA2 will use the truncation operator to remove the redundant solutions one by one, until $|A^{\prime}|=\bar{\mu}$. Let $D_{\bm{z}}=\{\bm{y}\in A^{\prime}\mid\bm{f}(\bm{z})=\bm{f}(\bm{y})\}$. Since $\bar{\mu}$ is at least $|\bm{f}(S)|$ and $|S|>\bar{\mu}$, there exists $\bm{z}\in A^{\prime}$ such that $|D_{\bm{z}}|>1$. Note that $\sigma_{\bm{x}}^{1}>0$ if $|D_{\bm{x}}|=1$, and $\sigma_{\bm{x}}^{1}=0$ if $|D_{\bm{x}}|>1$. Hence, if $|D_{\bm{x}}|=1$, then $\sigma_{\bm{x}}^{1}>0=\sigma_{\bm{z}}^{1}$, implying that $\bm{z}\leq_{d}\bm{x}$. Thus, $\bm{z}$ will be removed before $\bm{x}$, which implies that the non-duplicated solutions (regarding the objective vectors) will not be removed from $A^{\prime}$. Therefore, the lemma holds. ∎
@@ 4.2 Running Time of SPEA2 on Benchmark Problems
For SPEA2 solving $m$OneMinMax, if the size $\bar{\mu}$ of the archive is at least $(2n/m+1)^{m/2}$, uniform selection is employed to select parent solutions, and bit-wise mutation or one-bit mutation is employed to generate offspring solutions, the expected number of fitness evaluations for finding the Pareto front is $O(\mu n\cdot\min\{m\log n,n\})$.
@@
For SPEA2 solving $m$LeadingOnesTrailingZeroes, if the size $\bar{\mu}$ of the archive is at least $(2n/m+1)^{m-1}$, uniform selection is employed to select parent solutions, and bit-wise mutation or one-bit mutation is employed to generate offspring solutions, the expected number of fitness evaluations for finding the Pareto front is $O(\mu n^{2})$.
@@
For SPEA2 solving $m$OneJumpZeroJump, if the size $\bar{\mu}$ of the archive is at least $(2n/m-2k+3)^{m/2}$, uniform selection is employed to select parent solutions, and bit-wise mutation is employed to generate offspring solutions, the expected number of fitness evaluations for finding the Pareto front is $O(\mu n^{k}\cdot\min\{mn$, $3^{m/2}\})$.
@@
By Lemma 3 in [48], if $k\leq n^{\prime}/2$, the maximum cardinality of the set of non-dominated solutions having different objective vectors is at most $(2n/m+1)^{m/2}$ for $m$OneJump-ZeroJump, implying $|\bm{f}(S)|\leq(2n/m-2k+3)^{m/2}$. According to Lemma 4.1 and Theorem 3.3, if the archive size $\bar{\mu}\geq(2n/m-2k+3)^{m/2}$, the whole Pareto front can be found in $O(\mu n^{k}\cdot\min\{mn$, $3^{m/2}\})$ fitness evaluations in expectation. ∎
@@ 5 Application to Other Algorithms
$m$OneMinMax: For the bi-objective OneMinMax problem (i.e., $m$OneMinMax with $m=2$), because the SEMO algorithm always maintains the non-dominated set, the expected number of fitness evaluations to find the entire Pareto front is $2\mu n\log n=O(n^{2}\log n)$ by Theorem 3.1, where the equality holds because the population size is at most the Pareto front size $n+1$. This aligns with Giel and Lehre’s result [27]. Zheng et al. [46] proved that for a population size $\mu$ exceeding $4$ times the Pareto front size (i.e., $\mu\geq 4(n+1)$), NSGA-II preserves the non-dominated set on OneMinMax. By Theroem 3.1, the expected number of fitness evaluations to find the Pareto front is $O(\mu n\log n)=O(n^{2}\log n)$, aligning with their result. Similarly, Bian et al. [6] and Nguyen et al. [35] proved that when the population size is greater than the Pareto front size (i.e., $n+1$), both SMS-EMOA and $(\mu+1)$SIBEA preserve the non-dominated set on OneMinMax, resulting in an expected number of fitness evaluations of $O(\mu n\log n)$, consistent with Theorem 3.1. For $m$OneMinMax with large constant $m$, Andre et al. [36] proved that NSGA-III preserves the non-dominated set on $m$OneMinMax, if it employs a set of reference points $\mathcal{R}_{p}$ with $p\geq 4n\sqrt{m}$. With a population size of $\mu\geq(2n/m+1)^{m/2}$, the expected number of fitness evaluations is $O(\mu n\log n)$, which is consistent with the result derived from Theorem 3.1.
@@
$m$LeadingOnesTrailingZeroes: For the bi-objective LeadingOnesTrailingZeroes problem (i.e., $m$LeadingOnesTrailingZeroes with $m=2$), Laumanns [30] and Giel [26] proved that the expected number of fitness evaluations of SEMO and GSEMO for finding the Pareto front is $O(n^{3})$. Brockhoff [7] proved that when the population size is greater than the Pareto front size (i.e., $n+1$), $(\mu+1)$SIBEA preserves the non-dominated set on LeadingOnesTrailingZeroes, the Pareto front can be found in $O(\mu n^{2})$ number of fitness evaluations in expectation. Zheng et al. [46] proved that for the population size at least $4(n+1)$, NSGA-II preserves the non-dominated set on LeadingOnesTrailingZeroes and the expected number of fitness evaluations is also $O(\mu n^{2})=O(n^{3})$. These results are all consistent with Theorem 3.2. For $m\geq 4$, Laumanns [30] proved that the expected running time of SEMO is $O(n^{m+1})$, which is consistent with the result in Theorem 3.2, because the maximal population size during the optimization procedure is at most $(2n/m+1)^{m-1}\leq n^{m-1}$. Andre et al. [36] proved that NSGA-III preserves the non-dominated set on $m$LeadingOnesTrailingZeroes, if it employs a set of reference points $\mathcal{R}_{p}$ with $p\geq 4n\sqrt{m}$. With a population size of $\mu\geq(2n/m+1)^{m-1}$, the expected number of fitness evaluations is $O(\mu n^{2})$, which is consistent with Theorem 3.2.
@@
$m$OneJumpZeroJump: For the bi-objective OneJumpZeroJump problem (i.e., $m$-OneJumpZeroJump with $m=2$), Doerr et al. [22] proved that the expected number of fitness evaluations for GSEMO finding the Pareto front is $O((n-2k)n^{k})$, which is consistent with Theorem 3.3, because the maximal population size during the optimization procedure is at most $(n-2k+3)$. When the population size is at least $4(n-2k+3)$ for NSGA-II [19] and $(n-2k+3)$ for SMS-EMOA [6], both algorithms preserve the non-dominated set on $m$-OneJumpZeroJump and can find the Pareto front within $O(\mu n^{k})$ number of fitness evaluations in expectation, also consistent with Theorem 3.3. For $m\leq\log n$, Zheng and Doerr [48] proved that when the population size is at least $(2n/m-2k+3)^{m/2}$, SMS-EMOA preserves the non-dominated set and can find the Pareto front within $O(M\mu n^{k})$ number of fitness evaluations in expectation, where $M=(2n/m-2k+3)^{m/2}$, while Theorem 3.3 shows that any MOEA that can preserve the non-dominated set can find the Pareto front in $O(3^{m/2}\mu n^{k})$ expected number of fitness evaluations, implying a tighter bound.
@@ 6 Conclusion and Discussion
In this paper, we give general theorems for analyzing the expected running time of MOEAs on the three common benchmark problemes, $m$OneMinMax, $m$LeadingOnes-Trailingzeroes and $m$OneJumpZeroJump. These theorems show that if the population of MOEAs preserves the objective vectors of non-dominated solutions, their running time can be upper bounded. We apply them to derive the expected running time of SPEA2 for the first time, and also to analyze other MOEAs, deriving results consistent with previously known ones. We hope these theorems can be useful for theoretical analysis of MOEAs in the future. The theorems also suggest that when the population size is large enough, the analytical behavior of MOEAs may tend to be similar. Thus, it would be interesting to analyze the approximation performance of MOEAs with small populations, to better understand their differences.
In a parallel theoretical work, Wietheger and Doerr [41] proved near-tight running time guarantees for GSEMO on $m$OneMinMax, $m$CountingOnesCountingzeroes, $m$LeadingOnesTrailingZeroes, and $m$OneJumpZeroJump, and also transferred similar bounds to SEMO, SMS-EMOA and NSGA-III. Note that their proof techniques are superior and the bounds are tighter, as they not only considered the probability $p$ of not finding a Pareto front point within a given number $T$ of iterations, but also refined this process by considering each block individually. Specifically, they considered the probability $p^{\prime}$ of not achieving the optimization goal for each block within the given time. Using the union bound, leads to $p\leq m^{\prime}p^{\prime}$, where $m^{\prime}$ denotes the number of blocks. Let $M$ represent the size of the Pareto front. Then, the probability of finding the complete Pareto front within $T$ iterations is at least $1-Mp\geq 1-Mm^{\prime}p^{\prime}$. Their analysis of the optimization of blocks using the union bound allowed them to obtain tighter time bounds.

### v2

@@ 4.1 Large Archive Preserves Non-dominated Solutions
Throughout the process of SPEA2, the next population $P$ consists of the mutated individuals from the archive $A$. That is, the archive $A$ in SPEA2 actually plays the role of population, while the population $P$ acts as offspring population. We now show that when the archive size $\bar{\mu}$ is always no less than the maximum cardinality of the set of non-dominated solutions having different objective vectors for any $m$-objective function $\bm{f}$, SPEA2 will preserve the non-dominated set on $\bm{f}$.
@@
For SPEA2 solving $\bm{f}$, let $A^{\prime}$ denote the non-dominated set of solutions in the union of the population $P$ and archive $A$. If the size $\bar{\mu}$ of archive is always no less than $|\bm{f}(A^{\prime})|$, then SPEA2 preserves the non-dominated set.
@@
Consider that all the non-dominated solutions in $P\cup A$ have been added to $A^{\prime}$ (line 4 of Algorithm 1). If the size of $A^{\prime}$ is at most $\bar{\mu}$, then all the non-dominated solutions will survive to the next generation. Thus, the lemma holds. If the size of $A^{\prime}$ is larger than $\bar{\mu}$ (i.e., $|A^{\prime}|>\bar{\mu}$), SPEA2 will use the truncation operator to remove the redundant solutions one by one, until $|A^{\prime}|=\bar{\mu}$. Let $D_{\bm{z}}=\{\bm{y}\in A^{\prime}\mid\bm{f}(\bm{z})=\bm{f}(\bm{y})\}$. Since $\bar{\mu}$ is at least $|\bm{f}(A^{\prime})|$ and $|A^{\prime}|>\bar{\mu}$, there exists $\bm{z}\in A^{\prime}$ such that $|D_{\bm{z}}|>1$. Note that $\sigma_{\bm{x}}^{1}>0$ if $|D_{\bm{x}}|=1$, and $\sigma_{\bm{x}}^{1}=0$ if $|D_{\bm{x}}|>1$. Hence, if $|D_{\bm{x}}|=1$, then $\sigma_{\bm{x}}^{1}>0=\sigma_{\bm{z}}^{1}$, implying that $\bm{z}\leq_{d}\bm{x}$. Thus, $\bm{z}$ will be removed before $\bm{x}$, which implies that the non-duplicated solutions (regarding the objective vectors) will not be removed from $A^{\prime}$. Therefore, the lemma holds. ∎
@@ 4.2 Running Time of SPEA2 on Benchmark Problems
For SPEA2 solving $m$OneMinMax, if the size $\bar{\mu}$ of the archive is at least $(2n/m+1)^{m/2}$, uniform selection is employed to select parent solutions, and bit-wise mutation or one-bit mutation is employed to generate offspring solutions, the expected number of fitness evaluations for finding the Pareto front is $O(\bar{\mu}n\cdot\min\{m\log n,n\})$.
@@
For SPEA2 solving $m$LeadingOnesTrailingZeroes, if the size $\bar{\mu}$ of the archive is at least $(2n/m+1)^{m-1}$, uniform selection is employed to select parent solutions, and bit-wise mutation or one-bit mutation is employed to generate offspring solutions, the expected number of fitness evaluations for finding the Pareto front is $O(\bar{\mu}n^{2})$.
@@
For SPEA2 solving $m$OneJumpZeroJump, if the size $\bar{\mu}$ of the archive is at least $(2n/m-2k+3)^{m/2}$, uniform selection is employed to select parent solutions, and bit-wise mutation is employed to generate offspring solutions, the expected number of fitness evaluations for finding the Pareto front is $O(\bar{\mu}n^{k}\cdot\min\{mn$, $3^{m/2}\})$.
@@
By Lemma 3 in [48], if $k\leq n^{\prime}/2$, the maximum cardinality of the set of non-dominated solutions having different objective vectors is at most $(2n/m-2k+3)^{m/2}$ for $m$OneJump-ZeroJump, implying $|\bm{f}(S)|\leq(2n/m-2k+3)^{m/2}$. According to Lemma 4.1 and Theorem 3.3, if the archive size $\bar{\mu}\geq(2n/m-2k+3)^{m/2}$, the whole Pareto front can be found in $O(\bar{\mu}n^{k}\cdot\min\{mn$, $3^{m/2}\})$ fitness evaluations in expectation. ∎
@@ 5 Application to Other Algorithms
$m$OneMinMax: For the bi-objective OneMinMax problem (i.e., $m$OneMinMax with $m=2$), because the SEMO algorithm always maintains the non-dominated set, the expected number of fitness evaluations to find the entire Pareto front is $O(\mu n\log n)$ by Theorem 3.1, where the equality holds because the population size is at most the Pareto front size $n+1$. This aligns with Giel and Lehre’s result [27]. Zheng and Doerr [46] proved that for a population size $\mu$ exceeding $4$ times the Pareto front size (i.e., $\mu\geq 4(n+1)$), NSGA-II preserves the non-dominated set on OneMinMax. Then, by Theroem 3.1, the expected number of fitness evaluations to find the Pareto front is $O(\mu n\log n)$, aligning with their result. Similarly, Bian et al. [6] and Nguyen et al. [35] proved that when the population size is greater than the Pareto front size (i.e., $n+1$), both SMS-EMOA and $(\mu+1)$SIBEA preserve the non-dominated set on OneMinMax, resulting in an expected number of fitness evaluations of $O(\mu n\log n)$, consistent with Theorem 3.1. For $m$OneMinMax with large constant $m$, Andre et al. [36] proved that NSGA-III preserves the non-dominated set on $m$OneMinMax, if it employs a set of reference points $\mathcal{R}_{p}$ with $p\geq 4n\sqrt{m}$, and with a population size of $\mu\geq(2n/m+1)^{m/2}$. Then, they derived that the expected number of fitness evaluations is $O(\mu n\log n)$, which is consistent with the result derived from Theorem 3.1.
@@
$m$LeadingOnesTrailingZeroes: For the bi-objective LeadingOnesTrailingZeroes problem (i.e., $m$LeadingOnesTrailingZeroes with $m=2$), Laumanns et al. [30] and Giel [26] proved that the expected number of fitness evaluations of SEMO and GSEMO for finding the Pareto front is $O(n^{3})$. Brockhoff [7] proved that when the population size is greater than the Pareto front size (i.e., $n+1$), $(\mu+1)$SIBEA preserves the non-dominated set on LeadingOnesTrailingZeroes, then the Pareto front can be found in $O(\mu n^{2})$ number of fitness evaluations in expectation. Zheng and Doerr [46] proved that for the population size at least $4(n+1)$, NSGA-II preserves the non-dominated set on LeadingOnesTrailingZeroes and the expected number of fitness evaluations is also $O(\mu n^{2})$. These results are all consistent with Theorem 3.2. For $m\geq 4$, Laumanns et al. [30] proved that the expected running time of SEMO is $O(n^{m+1})$, which is consistent with the result in Theorem 3.2, because the maximal population size during the optimization procedure is at most $(2n/m+1)^{m-1}\leq n^{m-1}$. Andre et al. [36] proved that NSGA-III preserves the non-dominated set on $m$LeadingOnesTrailingZeroes, if it employs a set of reference points $\mathcal{R}_{p}$ with $p\geq 4n\sqrt{m}$, and with a population size of $\mu\geq(2n/m+1)^{m-1}$. Then, they derived that the expected number of fitness evaluations is $O(\mu n^{2})$, which is consistent with Theorem 3.2.
@@
$m$OneJumpZeroJump: For the bi-objective OneJumpZeroJump problem (i.e., $m$-OneJumpZeroJump with $m=2$), Doerr et al. [22] proved that the expected number of fitness evaluations for GSEMO finding the Pareto front is $O((n-2k)n^{k})$, which is consistent with Theorem 3.3, because the maximal population size during the optimization procedure is at most $(n-2k+3)$. When the population size is at least $4(n-2k+3)$ for NSGA-II [19] and $(n-2k+3)$ for SMS-EMOA [6], both algorithms preserve the non-dominated set on OneJumpZeroJump and can find the Pareto front within $O(\mu n^{k})$ number of fitness evaluations in expectation, also consistent with Theorem 3.3. For $m\leq\log n$, Zheng and Doerr [48] proved that when the population size is at least $(2n/m-2k+3)^{m/2}$, SMS-EMOA preserves the non-dominated set and can find the Pareto front within $O(M\mu n^{k})$ number of fitness evaluations in expectation, where $M=(2n/m-2k+3)^{m/2}$, while Theorem 3.3 shows that any MOEA that can preserve the non-dominated set can find the Pareto front in $O(3^{m/2}\mu n^{k})$ expected number of fitness evaluations, implying a tighter bound.
@@ 6 Conclusion and Discussion
In this paper, we give general theorems for analyzing the expected running time of MOEAs on the three common benchmark problems, $m$OneMinMax, $m$LeadingOnes-Trailingzeroes and $m$OneJumpZeroJump. These theorems show that if the population of MOEAs preserves the objective vectors of non-dominated solutions, their running time can be upper bounded. We apply them to derive the expected running time of SPEA2 for the first time, and also to analyze other MOEAs, deriving results consistent with previously known ones. We hope these theorems can be useful for theoretical analysis of MOEAs in the future. The theorems also suggest that when the population size is large enough, the analytical behavior of MOEAs may tend to be similar. Thus, it would be interesting to analyze the approximation performance of MOEAs with small populations, to better understand their differences.
In a parallel theoretical work, Wietheger and Doerr [41] proved near-tight running time guarantees for GSEMO on $m$OneMinMax, $m$CountingOnesCountingzeroes, $m$LeadingOnesTrailingZeroes, and $m$OneJumpZeroJump, and also transferred similar bounds to SEMO, SMS-EMOA and NSGA-III. Note that their proof techniques are superior and the bounds are tighter, as they not only considered the probability $p$ of not finding a Pareto front point within a given number $T$ of iterations, but also refined this process by considering each block individually. Specifically, they considered the probability $p^{\prime}$ of not achieving the optimization goal for each block within the given time. Using the union bound leads to $p\leq m^{\prime}p^{\prime}$, where $m^{\prime}$ denotes the number of blocks. Let $M$ represent the size of the Pareto front. Then, the probability of finding the complete Pareto front within $T$ iterations is at least $1-Mp\geq 1-Mm^{\prime}p^{\prime}$. Their analysis of the optimization of blocks using the union bound allows them to obtain tighter time bounds.